OpenStudy (lacris):

please help :( I will give you a shiny medal :D 2x-3[2x-(3-4x)] = ?? I need to simplify but I seem be getting a wrong answer. please give me detailed instructions, I would like to learn. Thank you :)

OpenStudy (kropot72):

The first step is to remove the inner brackets. 2x[2x - (3 - 4x)]=2x(2x -3 + 4x) Next simplify the expression inside the brackets, giving 2x(6x - 3) Now multiply out the brackets. Can you do that?

OpenStudy (lacris):

so it's 12x^2 - 6x? But I dont understand why we can remove the inner brackets? I thought we had to follow the rules and multiply everything out in the brackets first? :o

OpenStudy (lacris):

@kropot72 hello? :o

OpenStudy (kropot72):

In cases where there are bracketed expressions within other bracketed expressions, you need to remove the innermost brackets first. In this case the inner brackets were removed by multiplying each term inside the brackets by -1: -(3-4x) = -3 + 4x

OpenStudy (lacris):

ahh I see -(3-4x) was actually, -1(3-4x) ahh I see that

OpenStudy (kropot72):

Good to see that you can follow :)

OpenStudy (lacris):

so the answer is just 12x^2 - 6x then? Because the answer choices have: 8x-9 , -8x +3, 9-16x, and 12x^2-24x+9. Did I do something wrong?

OpenStudy (lacris):

@jhonyy9 can you help me? :D

OpenStudy (kropot72):

None of the answer choices that you posted relate to your question. You need to check everything.

OpenStudy (lacris):

@agent0smith I'm not sure if I'm doing this correctly

OpenStudy (lacris):

I keep getting 12x-9 for the answer..

OpenStudy (lacris):

is it the A^2+B^2 = (A+B)(A-B) law a part of this?

OpenStudy (agent0smith):

2x[2x-(3-4x)] = 12x^2 - 6x So either you wrote it wrong, or something. Screenshot the question maybe.

OpenStudy (lacris):

oh my gosh, I didn't even recognize that my keyboard didn't register the 2x-3 [ 2x-(3-4x)] .. I feel bad for giving the kropot a headache

OpenStudy (lacris):

oops.. but along the steps I do get 12x-18x-6x+9 when I try to simplify it and end up with 12x+9 :/ I'm not sure which part I'm simplifying wrong

OpenStudy (lacris):

yeah I will just open a new question.

OpenStudy (agent0smith):

Yeah just start another. Make sure the original problem is very clear.

OpenStudy (lacris):

yeah, it was my bad, oopsies D: